# Josephson traveling-wave parametric amplifier based on low-intrinsic-loss coplanar lumped-element waveguide

### v2

@@ Abstract
We present a Josephson traveling-wave parametric amplifier (JTWPA) based on a low-loss coplanar lumped-element waveguide architecture. By employing open-stub capacitors and Manhattan-pattern junctions, our device achieves an insertion loss below 1 dB up to 12 GHz. We introduce windowed sinusoidal modulation for phase matching, demonstrating that smooth impedance transitions effectively suppress intrinsic gain ripples. Using Tukey-windowed modulation with 8% impedance variation, we achieve 2023-dB gain over 5-GHz bandwidth under ideal matching conditions. In a more practical circuit having impedance mismatches, the device maintains 1720-dB gain over 4.8-GHz bandwidth with an added noise of 0.13 quanta above standard quantum limit at 20-dB gain and $-99$-dBm saturation power, while featuring zero to negative backward gain below the bandgap frequency.
@@ II.1 Open-stub transmission line structure
Our first objective is to obtain an underlying nonlinear transmission line with low insertion loss. We achieved this by building our device based on a coplanar waveguide (CPW), which is known for its low loss per unit length and ease of fabrication [34]. We introduced nonlinearity into the CPW by inserting Josephson junctions along its centerline to form a series junction array. To compensate for the increased series inductance due to the junctions and maintain a 50-$\Omega$ line impedance, additional shunt capacitance is required. To avoid using lossy dielectric materials, we employed open-stub shunt capacitors, where the low-loss silicon substrate and vacuum are the dominating dielectrics, as depicted in Fig. 1(b). The open-stub capacitors are constructed by extending part of the CPW centerline into both sides of the ground plane, added for each junction. A CP-JTWPA unit cell thus consists of a combination of an open-stub pair and a junction. These cells then chain up in series to form the underlying open-stub lines. The CP-JTWPA unit cell can be effectively modeled by an equivalent lumped-element circuit as shown in Fig. 1(c), consisting of the nonlinear inductance $L_{\mathrm{J}}$ and capacitance $C_{\mathrm{J}}$ of the series junction, and the shunt capacitance $C_{\mathrm{g}}$ of the open-stub pair. Additionally, we included two series inductors to take into account any stray inductance $L_{\mathrm{s}}$ due to the geometry.
To extract the parameters from the geometry, we perform finite-element simulation over a short chain of 40 unit cells, as simulating the full device ($>$2000 cells) is computationally prohibitive. We varied the length of the open stubs and compute the S-parameters for each length, followed by fitting them using the simulated results based on the lumped-element model using the ABCD-matrix method (see Appendix C). With the established scaling between stub length and shunt capacitance, we can achieve the desired impedance modulation by adjusting the stub length in each unit cell. In our implementation, we modulated both the junction area and stub length in a coordinated manner, that is, reducing the junction area (increasing $L_{\mathrm{J}}$) when the stub length is decreased (reducing $C_{\mathrm{g}}$) and vice versa. This correlated modulation allows an impedance variation without a large swing in stub lengths. Our Monte Carlo simulations (see Appendix E) reveal that this correlated modulation scheme preserves device performance for junction critical-current variations up to 8%, showing resilience comparable to capacitive-only modulation, despite the junctions no longer having the same critical current.
@@ II.2 Expanded-area Manhattan-type junctions
Building upon these design principles, we fabricated and characterized three CP-JTWPA samples in double-spiral layouts, housing up to 2400 unit cells each in a chip area of 5$\times$5 mm2, using the recipes described in Appendix G. We will refer to these as device A, B, and C throughout this text, each implementing a different window function (boxcar, Hann, and Tukey, respectively) on the sinusoidal impedance modulation (see Appendix A for detailed parameters).
A first achievement of our design is the low insertion loss, which we confirmed by measuring device B in a dilution refrigerator. By comparing the device transmission to a 50-$\Omega$ SubMiniature Version A (SMA) Female-to-Female (F–F) adapter using identical microwave cables sandwiched by microwave switches, we demonstrated insertion loss below 1 dB up to 12 GHz, including the loss from the sample package, as shown in Fig. 1(f). This low-loss performance addresses a persistent challenge in JTWPA development and establishes a foundation for amplifiers with noise performance approaching the standard quantum limit (SQL), as we will demonstrate in subsequent sections.
@@ III.1 Theoretical analysis
Figure 3: Comparison of transmission characteristics between device A (boxcar window, top) and device B (Hann window, bottom). Left panels [(a) and (b)] show both measured and simulated data from 412 GHz: measured and simulated normalized forward transmission $|S_{21}|^{2}$ without pump (transmission), measured and simulated normalized forward transmission $|S_{21}|^{2}$ with pump (gain), and simulated reflection $|S_{11}|^{2}$ (reflection). All simulations were performed assuming lossless transmission lines. Right panels [(c) and (d)] present enlarged views of the high-gain regions marked by red dotted-line boxes. Base transmission and reflection magnitudes are calculated using ABCD-matrix methods, while gain profiles are obtained through harmonic-balance simulations. The bandgap appears at 7.90 GHz and 8.13 GHz for devices A and B, respectively.
Conventional JTWPA designs typically implement uniform impedance modulation along the transmission line, which mathematically corresponds to applying a rectangular (boxcar) window to the spatial impedance profile. As shown in Fig. 2, this approach creates abrupt transitions at the JTWPA boundaries, leading to pronounced ripples in both transmission and reflection characteristics near the bandgap frequency. These transmission ripples emerge through a mechanism familiar in signal processing, where a finite-duration signal produces ripples in its frequency-domain spectrum due to windowing effects. Parametric gain in JTWPAs inherently amplifies any features present in the base transmission characteristics, so small ripples in transmission can lead to significant variations in the gain profile. These gain ripples pose practical limitations as they can cause frequency-dependent variations in signal-to-noise-ratio (SNR) improvement and complicate frequency-multiplexed readout. Moreover, the amplitude of ripples increases with the JTWPA gain, and their density increases with electrical length of the line, as they originate from the Fabry–Pérot-like interference due to imperfect impedance matching at the terminations [40]. In some previous work on JTWPAs based on periodic modulations, gain ripples were not as strong as predicted by simulations [11], which can be attributed to the damping effect from intrinsic transmission line losses. The achievement of low insertion loss in our CP-JTWPA devices removes this inadvertent damping mechanism, necessitating the suppression of gain ripples through deliberate design choices.
@@ III.2 Experimental validation
To validate our windowed modulation approach at higher modulation amplitudes, we characterized two CP-JTWPA variants, device A with a conventional boxcar window (5% impedance modulation) and device B with a Hann window (8% impedance modulation). Both devices were designed with a bandgap around $8$ $\mathrm{GHz}$ based on our open-stub-line architecture, with design parameters determined through ABCD-matrix analysis since the analytical coupled-mode theory becomes insufficient at high modulation amplitudes. We performed gain analysis using harmonic-balance simulations [41] to account for inter-cell reflections from both impedance mismatches and periodic modulations (see Appendix D for details). The devices were fabricated with similar processes and measured in dilution refrigerators over different cool-downs. In their microwave networks, they were sandwiched between cryogenic attenuators to minimize external impedance mismatches, allowing us to focus on their intrinsic characteristics. Given their low insertion loss ($<$$\mathrm{dB}$ up to 12 GHz, Fig. 1), in the following text, we will simply present the normalized parametric gain of our CP-JTWPAs by comparing pump-on and pump-off transmissions. For gain extraction near the bandgap, where pump-off transmission contains the stopband feature, we constructed the baseline by interpolating the transmission over the bandgap region from adjacent frequencies, which was then used to normalize the pump-on transmission to obtain the gain profile.
@@
Figure 4: Characterization of device C implementing a windowed periodic impedance modulation with a Tukey window with equal length in taper and maximum-amplitude region ($\alpha=0.5$). (a) The impedance modulation profile along the signal path. The modulation amplitude increases smoothly from zero at the input to 8% in the central uniform section, then decreases symmetrically to zero at the output. (b) Measured and simulated scattering parameters with 10-dB attenuators at both ports providing 50-$\Omega$ impedance matching. (c) Measured forward and backward scattering parameters without output attenuation (see Appendix F for the corresponding microwave network), representing a more realistic microwave environment with typical circuit impedance mismatches present in applications. The noise measurement was performed at 5.563 GHz (arrow) under this configuration.
@@
In Fig. 3, we present the scattering characteristics of CP-JTWPAs with box-windowed and Hann-windowed modulation profiles. Our initial measurements of the boxcar device used 3-dB attenuators at both ports, as the impact of impedance mismatches on JTWPA performance was not yet fully appreciated at that stage of development. Although this measurement configuration differs from later optimized setups, the observed ripple characteristics remain representative of the device’s intrinsic behavior due to their fundamental origin in the uniform modulation approach.
@@
The base transmission of the boxcar device exhibits ripples near the bandgap that match theoretical predictions for uniformly modulated transmission lines. Our simulation reveals that these ripples are accompanied by substantial reflection magnitudes in an oscillatory manner. These reflections create two critical challenges for JTWPA performance. First, the enhanced reflection near the bandgap complicates optimal pump placement. Although stronger phase correction requires placing the pump frequency closer to the bandgap, increased pump reflection in this region reduces effective pump power and can create standing waves that might promote unwanted reversed and reflected gain. In severe cases, these standing waves can prematurely saturate the Josephson junctions, preventing the JTWPA from achieving high gain. Second, the near-bandgap region where phase-matching conditions are better satisfied typically provides the highest gain but is also most affected by these ripples.
@@
When pumping the boxcar device, we observed a gain of $<$13 $\mathrm{dB}$ with significant gain ripples near the bandgap, agreeing well with the prediction by harmonic balance simulations. In particular, the frequency spacing between adjacent gain ripples closely matches that of the reflection ripples in the linear response, providing strong evidence that gain ripples originate directly from the underlying waveguide’s linear characteristics. These variations in gain can be seen in both measured and simulated data, and effectively limit the usable bandwidth for applications requiring uniform gain. As indicated in the enlarged view of the 57-GHz region, gain variations can exceed ±$\pm 5$ $\mathrm{dB}$ in the boxcar device. Our theoretical model in Eq. (B.32) further indicates that transmission ripples and consequently gain ripples become denser with increased JTWPA cell count, suggesting that scaling up boxcar-window JTWPAs to achieve higher gain would further increase the severity of this limitation.
@@
Next, we consider the transmission characteristics of the Hann device (lower panels of Fig. 3). Having learned the importance of minimizing external impedance mismatches, we measured this device with 6-dB attenuators to better isolate its intrinsic characteristics. While its larger bandgap width reflects the stronger modulation, it lacks the periodic oscillations in its base transmission near the bandgap that were characteristic of the boxcar device. The measured transmission shows only irregular fluctuations, likely arising from fabrication imperfections. Our simulations predict significantly reduced reflections near the bandgap for the Hann device with a sharp roll-off to $-20$ $\mathrm{dB}$, in contrast to the slow oscillatory decay observed in the boxcar device. This reduced reflection enables pump placement closer to the bandgap without the detrimental effect of pump reflection that the boxcar device might exhibit, allowing for stronger phase correction. The experimental results confirm this advantage through the achievement of phase-matched gain over a broad frequency range, delivering 1618 $\mathrm{dB}$ of gain over an instantaneous bandwidth of 3 GHz. Despite having a higher gain than the boxcar device, as suggested by the simulation, the gain profile of the Hann device exhibits significantly smoother characteristics, particularly near the bandgap where the strong ripples observed in the boxcar device are absent. The gain variations in the Hann device remain within ±$\pm 2$ $\mathrm{dB}$, validating windowed periodic modulation as an effective method to reduce gain ripples.
@@ IV.1 Gain characteristics
We implemented this approach in device C, with its impedance modulation profile shown in Fig. 4(a). To study its intrinsic transmission characteristics, we measured the device between 10-dB attenuators. The result is shown in Fig. 4(b). Simulations predicted minimal ripples in base transmission and faster reflection roll-off away from the bandgap compared to the boxcar device. By applying a pump slightly below the bandgap frequency, we achieved a gain of 2023 dB over a 5-GHz bandwidth. This gain significantly exceeded that of the Hann device overall, despite both devices shared similar designs except for their window functions. The Tukey device maintained excellent ripple suppression comparable to the Hann case, showing no strong ripples like those in the boxcar device even at higher gain.
@@
The measurements with attenuators demonstrated the intrinsic gain characteristics of CP-JTWPAs by neglecting the effects of background impedance mismatches. To study its behavior in a more realistic environment, we moved the Tukey device to a modified setup where we reduced the attenuation between the device and input circulator from 10 dB to 7 dB, and completely removed the 10-dB attenuator at the output port, thereby exposing the JTWPA directly to potential impedance mismatches from downstream components such as the output-side circulators (see Fig. 9). The measured forward and backward gain are shown in Fig. 4(c). We note that numerical fitting becomes impractical here due to insufficient knowledge of the frequency-dependent background impedance. The forward gain exhibits irregular ripples characteristic of impedance mismatches, with reduced overall amplitude compared to the impedance-matched case. This reduction stems from pump reflections generating standing waves, which can create local current amplitudes exceeding the junction critical current. This limitation forced us to reduce the input pump power, consequently lowering the achievable gain. Additionally, the gain shows faster roll-off away from the pump frequency, as the combined reflections from background impedance and bandgap necessitate placing the pump further from the bandgap, reducing phase-matching effectiveness. Nevertheless, the backward gain remains zero or negative across the useful bandwidth. The small negative backward transmission below the bandgap frequency suggested coupling of backward-propagating waves either to different frequencies or to the forward direction. While the precise mechanism requires further investigation, this coupling behavior could potentially be developed into an integrated approach for backward isolation in JTWPA design.
@@
Figure 5: Noise performance of the Tukey CP-JTWPA measured at 5.563 GHz. (a) Added noise versus gain. The added noise increases from $<$0.5 quanta at gain below 5 dB to 0.63 quanta at 21-dB gain, closely following the trend given by the SQL, before rising significantly at higher gain. (b) System noise temperature referred to the CP-JTWPA input. With the JTWPA off (orange), the system exhibits a baseline noise of approximately 6.3 K (23.6 quanta). When the JTWPA is activated (blue), the noise rapidly decreases to below 854 mK (3.2 quanta) within the first 10 dB of gain, then gradually approaches 320 mK (1.2 quanta) at 21-dB gain, before increasing at higher gain. Shaded regions represent statistical measurement uncertainties, with the added noise varying by $\pm$0.16 quanta at high gain. The system noise uncertainty ranges from $\pm$670 mK ($\pm$ 2.5 quanta) in the off state to $\pm$40 mK ($\pm$0.15 quanta) at high gain.

### v3

@@ Abstract
We present a Josephson traveling-wave parametric amplifier (JTWPA) based on a low-loss coplanar lumped-element waveguide architecture. By employing open-stub capacitors and Manhattan-pattern junctions, our device achieves an insertion loss below 1 dB up to 12 GHz. We introduce windowed sinusoidal modulation for phase matching, demonstrating that a smooth transition in the impedance-modulation strength effectively suppresses intrinsic gain ripples. Using Tukey-windowed modulation with 8% impedance variation, we achieve 20–23-dB gain over 5-GHz bandwidth under ideal matching conditions. In a more practical circuit having impedance mismatches, the device maintains 17–20-dB gain over 4.8-GHz bandwidth with an added noise of 0.18 quanta above standard quantum limit at 20-dB gain and 99-99-dBm saturation power, while featuring zero to negative backward gain below the band-gap frequency.
@@ II.1 Open-stub transmission line structure
Our first objective is to obtain an underlying nonlinear transmission line with low insertion loss. We achieved this by building our device based on a coplanar waveguide (CPW), which is known for its low loss per unit length and ease of fabrication [34]. We introduced nonlinearity into the CPW by inserting Josephson junctions along its centerline to form a series junction array. To compensate for the increased series inductance due to the junctions and maintain a 50-$\Omega$ line impedance, additional shunt capacitance is required. To avoid using lossy dielectric materials, we employed open-stub shunt capacitors, where the low-loss silicon substrate and vacuum are the dominating dielectrics, as depicted in Fig. 1(b). The open-stub capacitors are constructed by extending part of the CPW centerline into both sides of the ground plane, added for each junction. A CP-JTWPA unit cell thus consists of a combination of an open-stub pair and a junction. These cells then chain up in series to form the underlying open-stub lines. The CP-JTWPA unit cell can be effectively modeled by an equivalent lumped-element circuit as shown in Fig. 1(c), consisting of the nonlinear inductance $L_{\mathrm{J}}$ and capacitance $C_{\mathrm{J}}$ of the series junction, and the shunt capacitance $C_{\mathrm{g}}$ of the open-stub pair. Additionally, we included two series inductors to take into account any stray inductance $L_{\mathrm{s}}$ due to the geometry. We note that a similar open-stub approach for JTWPA implementation has been recently demonstrated [35] also as an attempt to achieve low insertion loss, though with a different phase-matching scheme rather than our windowed periodic-modulation approach.
To extract the parameters from the geometry, we perform finite-element simulations over a short chain of 40 unit cells, as simulating the full device (>>2000 cells) is computationally prohibitive. We varied the length of the open stubs and computed the S-parameters for each length, followed by fitting the results to the lumped-element model using the ABCD-matrix method (see Appendix C). With the established scaling between stub length and shunt capacitance, we can achieve the desired impedance modulation by adjusting the stub length in each unit cell.
In our implementation, we modulated both the junction area and stub length in a coordinated manner, in order to achieves the required 8% impedance modulation. That is, moving along the series junction array, we reduce the junction area (increasing $L_{\mathrm{J}}$) when the stub length is decreased (reducing $C_{\mathrm{g}}$) and vice versa. This correlated approach reduces the required stub-length variation from approximately 16% (if modulating capacitance alone) to 8%, as detailed in Appendix A, thus avoiding a large geometric variation in the design. We verified through Monte Carlo simulations (see Appendix F) that this correlated modulation scheme maintains resilience against junction critical-current variations comparable to capacitive-only modulation, despite the junction areas are no longer uniform by design.
@@ II.2 Expanded-area Manhattan-type junctions
Building upon these design principles, we fabricated and characterized three CP-JTWPA samples in double-spiral layouts, all housing roughly 2400 unit cells each in a chip area of $\times5$ mm2, using the recipes described in Appendix H. We will refer to these as device A, B, and C throughout this text, each implementing a different window function (boxcar, Hann, and Tukey, respectively) on the sinusoidal impedance modulation (see Appendix A for detailed parameters).
A first achievement of our design is the low insertion loss, which we confirmed by measuring device B in a dilution refrigerator. By comparing the device transmission to a 50-$\Omega$ SubMiniature Version A (SMA) Female-to-Female (F–F) adapter using identical microwave cables sandwiched by microwave switches, we demonstrated insertion loss below 1 dB up to 12 GHz, including the loss from the sample package, as shown in Fig. 1(f). Consistent low-loss transmissions are also observed for device A and C. This low-loss performance addresses a persistent challenge in JTWPA development and establishes a foundation for amplifiers with noise performance approaching the standard quantum limit (SQL), as we will demonstrate in subsequent sections.
@@ III.1 Theoretical analysis
Figure 3: Comparison of transmission characteristics between device A (boxcar window, top) and device B (Hann window, bottom). Left panels [(a) and (b)] show both measured and simulated data from 4–12 GHz: measured and simulated normalized forward transmission $|S_{21}|^{2}$ without pump (transmission), measured and simulated normalized forward transmission $|S_{21}|^{2}$ with pump (gain), and simulated reflection $|S_{11}|^{2}$ (reflection). All simulations were performed assuming lossless transmission lines. Right panels [(c) and (d)] present enlarged views of the high-gain regions marked by red dotted-line boxes. Base transmission and reflection magnitudes are calculated using ABCD-matrix methods, while gain profiles are obtained through harmonic-balance simulations. The band gap appears at 7.90 GHz and 8.13 GHz for devices A and B, respectively.
Conventional JTWPA designs typically implement uniform impedance modulation along the transmission line, which mathematically corresponds to applying a rectangular (boxcar) window to the spatial impedance profile. As shown in Fig. 2, this approach creates abrupt transitions at the JTWPA boundaries, leading to pronounced ripples in both transmission and reflection characteristics near the band-gap frequency. These transmission ripples emerge through a mechanism familiar in signal processing, where a finite-duration signal produces ripples in its frequency-domain spectrum due to windowing effects. Parametric gain in JTWPAs inherently amplifies any features present in the base transmission characteristics, so small ripples in transmission can lead to significant variations in the gain profile. These gain ripples pose practical limitations as they can cause frequency-dependent variations in signal-to-noise-ratio (SNR) improvement and complicate frequency-multiplexed readout. Moreover, the amplitude of ripples increases with the JTWPA gain, and their density increases with electrical length of the line, as they originate from the Fabry–Pérot-like interference due to impedance mismatches at the terminations between the periodic structure’s Bloch impedance and the external 50-$\Omega$ environment. [41, 42]. In some previous work on JTWPAs based on periodic modulations, gain ripples were not as strong as predicted by simulations [11], which can be attributed to the damping effect from intrinsic transmission line losses. The achievement of low insertion loss in our CP-JTWPA devices removes this inadvertent damping mechanism, necessitating the suppression of gain ripples through deliberate design choices.
@@ III.2 Experimental validation
To validate our windowed modulation approach at higher modulation amplitudes, we characterized two CP-JTWPA variants, device A with a conventional boxcar window (5% impedance modulation) and device B with a Hann window (8% impedance modulation). Both devices were designed with a band gap around 88 $\mathrm{GHz}$ based on our open-stub-line architecture, with design parameters determined through ABCD-matrix analysis since the analytical coupled-mode theory becomes insufficient at high modulation amplitudes. We performed gain analysis using harmonic-balance simulations [43] to account for inter-cell reflections from both impedance mismatches and periodic modulations (see Appendix D for details). The devices were fabricated with similar processes and measured in dilution refrigerators over different cool-downs. In their microwave networks, they were sandwiched between cryogenic attenuators to minimize external impedance mismatches, allowing us to focus on their intrinsic characteristics. Given their low insertion loss ($\mathrm{dB}$ up to 12 GHz, Fig. 1), in the following text, we will simply present the normalized parametric gain of our CP-JTWPAs by comparing pump-on and pump-off transmissions. For gain extraction near the band gap, where pump-off transmission contains the stopband feature, we constructed the baseline by interpolating the transmission over the band-gap region from adjacent frequencies, which was then used to normalize the pump-on transmission to obtain the gain profile.
@@
Figure 4: Characterization of device C implementing a windowed periodic impedance modulation with a Tukey window with equal length in taper and maximum-amplitude region ($\alpha=0.5$). (a) The impedance modulation profile along the signal path. The modulation amplitude increases smoothly from zero at the input to 8% in the central uniform section, then decreases symmetrically to zero at the output. (b) Measured and simulated scattering parameters with 10-dB attenuators at both ports providing 50-$\Omega$ impedance matching. (c) Measured forward and backward scattering parameters without output attenuation (see Appendix G for the corresponding microwave network), representing a more realistic microwave environment with typical circuit impedance mismatches present in applications. The noise measurement was performed at 5.563 GHz (arrow) under this configuration.
@@
In Fig. 3, we present the scattering characteristics of CP-JTWPAs with boxcar-windowed and Hann-windowed modulation profiles. Our initial measurements of the boxcar device used 3-dB attenuators at both ports, as the impact of impedance mismatches on JTWPA performance was not yet fully appreciated at that stage of development. We progressively increased the installed attenuation to 6 dB (Hann) and 10 dB (Tukey) as we observed higher gains from successive samples. While these different attenuation values might initially appear problematic, we demonstrate quantitatively in Appendix E that these measurement configurations do not bias the experimental results and that the observed characteristics represent the intrinsic device behavior.
@@
The baseline transmission of the boxcar device exhibits ripples near the band gap that match theoretical predictions for uniformly modulated transmission lines. Our simulations reveal that these ripples are accompanied by substantial reflection magnitudes exhibiting oscillatory behavior. These reflections create two critical challenges for JTWPA performance. First, enhanced reflection near the band-gap complicates optimal pump placement. Although stronger phase correction requires positioning the pump frequency closer to the band gap, increased pump reflection in this region reduces effective pump-power delivery and can generate standing waves with elevated local current amplitudes, potentially saturating certain junctions prematurely. Second, while near-band-gap regions typically provide better phase-matching conditions and the highest gain, they also exhibit the strongest gain ripples, making these desired frequency ranges unavailable for applications sensitive to gain fluctuations.
@@
When pumping the boxcar device, we observed a gain of <<13 $\mathrm{dB}$ with significant gain ripples near the band gap, agreeing well with the prediction by harmonic balance simulations. In particular, the frequency spacing between adjacent gain ripples closely matches that of the reflection ripples in the linear response, providing strong evidence that gain ripples originate directly from the underlying waveguide’s linear characteristics. These variations in gain can be seen in both measured and simulated data, and effectively limit the usable bandwidth for applications requiring uniform gain. As indicated in the enlarged view of the 5–7-GHz region, gain variations can exceed $\pm 5$ $\mathrm{dB}$ in the boxcar device. Our theoretical model in Eq. (B.32) further indicates that transmission ripples and consequently gain ripples become denser with increased JTWPA cell count, suggesting that scaling up boxcar-window JTWPAs to achieve higher gain would further increase the severity of this limitation.
@@
Next, we consider the transmission characteristics of the Hann device (lower panels of Fig. 3) by measuring it sandwiched between two 6-dB attenuators. While its larger band-gap width reflects the stronger modulation, it lacks the periodic oscillations in its base transmission near the band gap that were characteristic of the boxcar device. The measured transmission shows only irregular fluctuations, likely arising from fabrication imperfections. Our simulations predict significantly reduced reflections near the band gap for the Hann device with a sharp roll-off to 20-20 $\mathrm{dB}$, in contrast to the slow oscillatory decay observed in the boxcar device. This reduced reflection enables pump placement closer to the band gap without the detrimental effect of pump reflection that the boxcar device might exhibit, allowing for stronger phase correction. The experimental results confirm this advantage through the achievement of phase-matched gain over a broad frequency range, delivering 16–18 $\mathrm{dB}$ of gain over an instantaneous bandwidth of 3 GHz. Despite having a higher gain than the boxcar device, as suggested by the simulation, the gain profile of the Hann device exhibits significantly smoother characteristics, particularly near the band gap where the strong ripples observed in the boxcar device are absent. The gain variations in the Hann device remain within $\pm 2$ $\mathrm{dB}$, validating windowed periodic modulation as an effective method to reduce gain ripples.
@@ IV.1 Gain characteristics
We implemented this approach in device C, with its impedance modulation profile shown in Fig. 4(a). To study its intrinsic transmission characteristics, we measured the device between 10-dB attenuators. The result is shown in Fig. 4(b). Simulations predicted minimal ripples in base transmission and faster reflection roll-off away from the band gap compared to the boxcar device. By applying a pump slightly below the band-gap frequency, we achieved a gain of 20–23 dB over a 5-GHz bandwidth. This gain significantly exceeded that of the Hann device overall, despite both devices shared similar designs except for their window functions. The Tukey device maintained excellent ripple suppression comparable to the Hann case, showing no strong ripples like those in the boxcar device even at higher gain.
@@
The measurements with attenuators demonstrated the intrinsic gain characteristics of CP-JTWPAs by neglecting the effects of background impedance mismatches. To study its behavior in a more realistic environment, we moved the Tukey device to a modified setup where we reduced the attenuation between the device and input circulator from 10 dB to 7 dB, and completely removed the 10-dB attenuator at the output port, thereby exposing the JTWPA directly to potential impedance mismatches from downstream components such as the output-side circulators (see Fig. 10). The measured forward and backward gain are shown in Fig. 4(c). We note that numerical fitting becomes impractical here due to insufficient knowledge of the frequency-dependent background impedance. The forward gain exhibits irregular ripples characteristic of impedance mismatches, with reduced overall amplitude compared to the impedance-matched case. This reduction stems from pump reflections generating standing waves, which can create local current amplitudes exceeding the junction critical current. This limitation forced us to reduce the input pump power, consequently lowering the achievable gain. Additionally, the gain shows faster roll-off away from the pump frequency, as the combined reflections from background impedance and band gap necessitate placing the pump further from the band-gap, reducing phase-matching effectiveness. Nevertheless, the backward gain remains zero or negative across the useful bandwidth. The small negative backward transmission below the band-gap frequency suggested coupling of backward-propagating waves either to different frequencies or to the forward direction. While the precise mechanism requires further investigation, this coupling behavior could potentially be developed into an integrated approach for backward isolation in JTWPA design.
@@
Figure 5: Noise performance of the Tukey CP-JTWPA measured at 5.563 GHz. (a) Added noise versus gain. The added noise increases from <<0.5 quanta at gain below 5 dB to 0.68 quanta at 21-dB gain, closely following the trend given by the SQL, before rising significantly at higher gain. (b) System noise temperature referred to the CP-JTWPA input. With the JTWPA off (orange), the system exhibits a baseline noise of approximately 7.0 K (26.2 quanta). When the JTWPA is activated (blue), the noise rapidly decreases to below 1.06 K (3.96 quanta) within the first 10 dB of gain, then gradually approaches 382 mK (1.43 quanta) at 21-dB gain, before increasing at higher gain. Shaded regions represent statistical measurement uncertainties, with the added noise varying by $\pm$0.31 quanta at high gain. The system noise uncertainty ranges from $\pm$1.72 K ($\pm$ 6.42 quanta) in the off state to $\pm$90 mK ($\pm$0.34 quanta) at high gain.